Maria averaged 212 miles on the first day of her trip. She averaged only 104 miles on the last six days of her trip. What was the overall average number of miles driven per day?

Answer by jrfrunner(365)

the overall average is really a weighted average
--
weighted average = total miles driven/total days driven = (212+104)/7=45.14
==
alternatively, you could figure the miles each day and multiply by the proportion represented by each day.
--
for example, day 1 miles=212, day 2 miles =104/6=17.33, day 3 miles=17.33 etc
each day represents 1/7 of the trip.
--
weighted average= 212*(1/7)+17.33*(1/7)+...17.33*(1/7)=212*(1/7)+17.33*(6/7)=45.14